OpenStudy (anonymous):

Write an equation in point slope form given: slope, m, = 2/7 and point (4, -1). A. y + 1 = -(2/7)(x – 4) B. y + 1 = (2/7)(x – 4) C. y – 1 = (2/7)(x – 4) D. y – 1 = -(2/7)(x – 4)

OpenStudy (anonymous):

only the middle two have the correct slope, so ignore 1 and 4

OpenStudy (anonymous):

you write \[y-y_1=m(x-x_1)\] puttin \(m=\frac{2}{7}, x_1=4, y_1=-1\)

OpenStudy (anonymous):

thank you! (: